OpenStudy (anonymous):

Which of the following is a requirement for a trinomial to be factored as a perfect square? A. The constant term must be a perfect square. B. The constant term must be a perfect cube. C. The constant term must be twice the product of the square roots of the other two terms. D. The constant term must be half the product of the square roots of the other two terms.

OpenStudy (anonymous):

well the answer would be c For example 4x^2 + 12xy + 9y^2 = (2x + 3y)^2

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm ax^2+bx+c }\) \(\large\color{midnightblue}{ \rm x^2+\frac{b}{a}x+\frac{c}{a} }\) \(\large\color{red}{ \rm (\frac{b}{2a})^2=\frac{c}{a} }\)

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

it is hard to understand what I am saying...

OpenStudy (anonymous):

yah it is

OpenStudy (solomonzelman):

I like numbers better than words when answering a MATH question.

OpenStudy (anonymous):

oh ok